Question 1:Compare ( Interpretation) the means of each pair of groups using the Bonferroni method.

Question 2: Compare ( Interpretation)the means of each pair of groups using the LSD method and SNK method.

Post Hoc Tests Multiple Comparisons Dependent Variable: ratioFevFvc Mean Difference ( J) 95% Confidence Interval Lower Bound 2.1907 3.8164 18.0193 4.7993 -22.6236 -11.0293 1788 1.4259 -20.0312 6.8112 -25.0141 13.0412 Upper Bound 18.0193 22.6236 2.1907 11.0293 3.8164 4.7993 20.0312 25.0141 .1788 13.0412 1.4259 6.8112 Groups Groups Group A Std. Error 3.78126 13.22000 4.49281 10.10500 3.78126 3.11500 3.78126 13.22000 4.49281 3.11500 3.78126 10.10500 3.78126 13.22000 4.49281 10.10500 3.78126 3.11500 3.78126 13.22000 4.49281 3.11500 3.78126 Sig 10.10500 015 008 015 420 008 420 045 025 045 1.000 025 1.000 LSD Group B Group C Group A Group C Group A Group B Group B Group C Group A Group C Group A Group B Group C Bonferroni GroupA Group B Group C *. The mean difference is significant at the 0.05 levelHomogeneous Subsets ratioFevFvc Subset for alpha 0.05 Groups Student-Newman-Keulsa Group C ,b 5.4600 8.5750 Group B Group A Sig 12 18.6800 449 1.000 Means for groups in homogeneous subsets are displayed. a. Uses Harmonic Mean Sample Size 6.207 b. The group sizes are unequal. The harmonic mean of the group sizes is used. Type l error levels are not guaranteed
Answer #1

Bonferroni method -
Mean difference between Group A and Group B = 10.105 with significance = 0.045.
As, the significance is less than 0.05, Mean difference between Group A and Group B is statistically significant.
That is, there is significant evidence that the mean difference between Group A and Group B is not 0.

Mean difference between Group A and Group C = 13.22 with significance = 0.025.
As, the significance is less than 0.05, Mean difference between Group A and Group C is statistically significant.
That is, there is significant evidence that the mean difference between Group A and Group C is not 0.

Mean difference between Group B and Group C = 3.115 with significance = 1.00.
As, the significance is greater than 0.05, Mean difference between Group B and Group C is not statistically significant.
That is, there is no significant evidence that the mean difference between Group B and Group C is not 0.


LSD method -
Mean difference between Group A and Group B = 10.105 with significance = 0.015.
As, the significance is less than 0.05, Mean difference between Group A and Group B is statistically significant.
That is, there is significant evidence that the mean difference between Group A and Group B is not 0.

Mean difference between Group A and Group C = 13.22 with significance = 0.008.
As, the significance is less than 0.05, Mean difference between Group A and Group C is statistically significant.
That is, there is significant evidence that the mean difference between Group A and Group C is not 0.

Mean difference between Group B and Group C = 3.115 with significance = 0.42.
As, the significance is greater than 0.05, Mean difference between Group B and Group C is not statistically significant.
That is, there is no significant evidence that the mean difference between Group B and Group C is not 0.

SNK Method -
Mean difference between Group A and Group B = 18.68 - 8.575 = 10.105
Mean difference between Group A and Group C = 18.68 - 5.46 = 13.22
Mean difference between Group C and Group C = 8.575 - 5.46 = 3.115
Group B and C lies in one group. Thus, Mean difference between Group A and Group B, C are statistically significant.